OpenStudy (vera_ewing):

Choose a chemical reaction involving two or more reactants. For this reaction, how do you determine which reactant or reactants are limiting? Explain.

OpenStudy (vera_ewing):

@Abmon98 help please!

OpenStudy (abmon98):

Where is the reaction given?

OpenStudy (vera_ewing):

i wasn't given one....i think we have to make one up

OpenStudy (dtan5457):

@vera_ewing Need an example?

OpenStudy (dtan5457):

Here. If C3H8+5O2>>3CO2+4H2O has 1.5 moles of C3H8 and 15 moles of O2 reacting... then you need to make two ratios, a problem ratio, and a equation ratio the problem ratio is the statement of (1.5 moles of C3H8 and 15 moles of O2 reacting) so the problem ratio=1.5/15=0.10 the equation ratio is what's actually on the equation, in this case there is 1 mole of C3H8 and 5 moles of O2. problem ratio=1/5=0.2 now the important part is...if the mole ratio in the problem is > than then the equation, then the numerator is excess, and the denominator is limiting. if the problem ratio < equation ratio, then the denominator is excess, and the numerator is limiting.